40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Engineering group Lab elog  Not logged in ELOG logo
Message ID: 225     Entry time: Fri Dec 13 10:57:23 2019
Author: Marie K. 
Type: Progress 
Category: BS BRDs 
Subject: Survey of BRD_v5 drifts on BS (Fall 2019) 

Attached are the measurements of the bounce roll transfer function of the dummy BS with BRD5 and BRD6 (v5 blades) fastened to the pum.

BRD5_v5 was built on July 16 (elog 165) and BRD6_v5 was built on August 21 (elog 198) after baking (elog 204). It was found that BRD5_v5 stopped drifting at the end of August (elog 204). They were attached on the suspension on August 28th after tuning (elog 202). The date of the measurements is recalled in the table attached, with the latests ones from December 4th, 2019. There are some measurements missing from early November unfortunately (issues with the spectrum analyzer saving data/USB key problems).

Nevertheless, we can clearly see that the BS resonances are still damped over the time period of 3 months, meaning that BRDs are still well in tune with the BS resonances. It means that they stopped drifting or that the drift is now negligeable.

Note: the measurements 206, 207, 208, 212 were taken with a lower amplitude of excitation (250 mV instead of 1V) to check if a high amplitude excitation was not introducing non-linear effects. It might be that the coherence is lower for these measurements and there are therefore less reliable.

 

Attachment 1: BRD5&6_Fall_2019_survey_bounce.png  263 kB  | Hide | Hide all
BRD5&6_Fall_2019_survey_bounce.png
Attachment 2: BRD5&6_Fall_2019_survey_roll.png  303 kB  | Hide | Hide all
BRD5&6_Fall_2019_survey_roll.png
Attachment 3: August2019_monitoring.xlsx  42 kB
ELOG V3.1.3-